# UrbanKGent: A Unified Large Language Model Agent Framework for Urban Knowledge Graph Construction

### v1

@@ 3.1.1. Task Definition
Relational Triplet Extraction (RTE). Given the unstructured texts, this task achieves joint extraction of entities and their relations (Zheng et al., 2021) which are in the form of a triplet ${\langle h,r,t\rangle}$. For instance, given the urban text sentence ”Columbia University is a private Ivy league research university in New York City.”, this task aims to identify two entities ”Columbia University” and ”New York City” and their relation ”locate-in”, described as triplet ¡Columbia University, locate-in, New York City¿.
Knowledge Graph Completion (KGC). Given a head entity $h$ and a tail entity $t$, this task is to predict the missing relation between them (Yao et al., 2023). For instance, given the head entity ”Columbia University” and the tail entity ”Empire State Building”, this task is to predict that their missing relation, e.g., ”disconnected”, described as triplet ¡Columbia University, disconnected, Empire State Building¿.
3.1.2. Problem Formulation
Given the urban unstructured text data, the desired output is an UrbanKG $\mathcal{G}$. In this paper, this problem is decomposed into two sequential subtasks: (1) Relational Triplet Extraction: the first task extracts relational triplet ${\langle h,r,t\rangle}$ from the urban text data. The output of RTE task is $\mathcal{G}_{1}$ $\mathcal{E}$, $\mathcal{R}_{1}$, $\mathcal{F}_{1}$), where $\mathcal{E}$ and $\mathcal{R}_{1}$is the set of extracted entities and relations, while $\mathcal{F}_{1}$ is the set of extracted triplets. (2) Knowledge Graph Completion: for the given head entity $h$ and tail entity $t$ in $\mathcal{G}_{1}$, the second task is to predict the geospatial relationship [^7] We follow GeoLM (Li et al., 2023c) to provide five RCC relationship (Mani et al., 2010) candidates: Disconnection (DC), external connection (EC), equality (EQ), partial overlap (PO), and tangential and non-tangential proper parts (IN). Details are in Appendix A. between them. The output of this task is $\mathcal{G}_{2}$ $\mathcal{E}$, $\mathcal{R}_{2}$, $\mathcal{F}_{2}$), where $\mathcal{R}_{2}$ and $\mathcal{F}_{2}$ is the set of completed relations and triplets. By sequentially completing the above two tasks, we can obtain the constructed UrbanKG $\mathcal{G}$ = ($\mathcal{E}$, $\mathcal{R}_{1}\cup\mathcal{R}_{2}$, $\mathcal{F}_{1}\cup\mathcal{F}_{2}$).
3.2. Quantitative Task Analysis
@@
Figure 3. Quantitative performance analysis of prompting GPT-4 for RTE and KGC tasks. The result is obtained by comparing 50 GPT-4’s outputs with the human’s annotation.
@@
Heterogenous Relationship Understanding. The ability to understand heterogeneous relationships is ubiquitous in distilling knowledge from the massive urban corpus. For example, the text description in Figure 2 illustrates a place from spatial location, temporal time, and functional aspects. Capturing these heterogeneous semantics is important for urban knowledge distillation. We perform quantitative analysis by randomly sampling 50 urban text data and then prompt GPT-4 to complete relational triplet extraction by providing only the basic task description. As shown in Figure 3(a), we find that the LLMs experience serious misjudgment (i.e., extract wrong triplets or miss the triplet) on urban spatial, temporal, and functional triplet extraction. This indicates the limited capacity of LLMs to understand heterogeneous relationships.
Geospatial Computing and Reasoning. Geospatial computing and reasoning techniques are widely used in many previous UrbanKG studies (Ning et al., 2023; Liu et al., 2023) for urban geospatial relation extraction. In recent works (Mooney et al., 2023; Manvi et al., 2023), the geospatial skills of LLMs have also been demonstrated to lack geospatial awareness and reasoning ability (Bhandari et al., 2023). To identify potential limitations, we quantitatively investigate how LLMs can perform on geospatial relation completion tasks. Specifically, we construct 100 head and tail entity pairs, covering five geospatial relations in the KGC task, and then prompt GPT-4 to predict with basic task description and geospatial relation candidates. As shown in Figure 3(b), we find that GPT-4 performs poorly on five geospatial relation completion. This further validates the disability of LLMs in geospatial computing and reasoning.
4. UrbanKGC Agent Construction
@@
4.1. Overview
The overall pipeline of the UrbanKGent framework is illustrated in Figure 4, which consists of three major components: (1) Knowledgeable Instruction Generation consists of the heterogeneity-aware and geospatial-infused instruction generation modules for aligning LLMs to UrbanKGC tasks. (2) Tool-augmented Iterative Trajectory Refinement proposes geospatial tool interface invocation and iterative self-refinement mechanisms to enhance and refine generated trajectory. (3) Hybrid Instruction Fine-tuning fine-tune LLMs based on the refined trajectories for cost-effectively completing diverse UrbanKGC tasks.
4.2. Knowledgeable Instruction Generation
@@
4.2.1. Heterogeneity-aware Instruction Generation for Relational Triplet Extraction.
@@
As discussed in Section 3, the urban text contains diverse heterogeneous relationships, thus we consider multiple views with both urban entity and relation definition for relational triplet extraction. In particular, we construct a multi-view instruction template for the urban relational triplet extraction, including spatial view, temporal view, and functional view. Each view is a multi-turn question-answer dialog (Wei et al., 2023) consisting of entity recognition, relation extraction, and triplet extraction module.
@@
For the spatial view, we devise a two-turn dialog to align LLMs for spatial triplet extraction. In the first turn, we inject spatial entity and relation definition into the instruction template to guide LLMs to understand spatial characteristics and then extract potential spatial entities (e.g., University) and relations types (e.g., locate-in). In the second turn, the extracted types are explicitly fed into the instruction template for spatial triplet extraction. Intuitively, the spatial view allocates dedicated urban knowledge for LLMs to extract urban spatial relationships. Similarly, we construct the temporal view and functional view for corresponding temporal and functional triplet extraction, independently. We provide the detailed instruction template in the Appendix B.
4.2.2. Geospatial-infused Instruction Generation for Knowledge Graph Completion.
@@
Despite heterogeneity-aware instruction enabling LLMs to extract urban triplets from various perspectives, the geospatial relationship between geospatial entities cannot be directly extracted. Therefore, we introduce a geospatial-infused instruction generation module to guide LLMs to complete missing geospatial relationships.
@@
First, we incorporate the geometry information (i.e., the latitude and longitude) of geo-entities into instruction, so that the LLMs can utilize these geospatial values for relation inference. Second, we add the geospatial relationship definition to the instruction to guide LLMs in understanding the geospatial relationship definition. Intuitively, LLMs can refer to this geospatial knowledge and make practical solutions for the knowledge graph completion task. We provide the detailed instruction template in the Appendix B.
4.3. Tool-augmented Iterative Trajectory Refinement
Then, we present the trajectory generation to further augment and refine the constructed UrbanKGC instructions.
4.3.1. Trajectory Generation
@@
With the initial UrbanKGC instructions constructed, the following step is to generate reasoning trajectories (Zeng et al., 2023), which will be used to fine-tune LLMs tailored to UrbanKGC task. Specifically, we follow FireAct (Chen et al., 2023) and use Chain-of-Thought (CoT) (Wei et al., 2022), a gradient-free technique, to prompt GPT-4 (i.e., add prompt trigger ”Let’s think step by step” at the end of RTE and KGC instructions template) to generate the reasoning trajectories for UrbanKGC tasks.
@@
The generated CoT trajectories could provide a step-by-step reasoning solution for UrbanKGC tasks. Nevertheless, the complex geospatial relationships cannot be easily extracted as discussed in Section 3 and recent geospatial reasoning works (Manvi et al., 2023; Mooney et al., 2023; Bhandari et al., 2023). Therefore, we introduce a tool invocation module to guide LLMs to invoke tailored external geospatial tools (Cai et al., 2023) to enhance their geospatial computing and reasoning capacity for UrbanKGC tasks.
Figure 4. An overview of UrbanKGent Construction.
4.3.2. Tool Invocation for Trajectory Augmentation
@@
Tool Invocation. First, we construct a geospatial reasoning toolkit (e.g., distance calculation, eight interfaces in total shown in Table 8) by prompting GPT-4 for self-programming. Then, we construct tailored prompts to guide LLMs to invoke these interfaces. Specifically, the prompt is concatenated with an illustrative description of the function of each geospatial tool and a task instruction (i.e., ”Which types of tool interface you need”). Intuitively, the external tool allocates necessary calculation results for LLMs to infer missing geospatial relation. The detailed toolkit description can be found in Appendix C.
Trajectory Deliberation. After manipulation with external tools, we prompt LLMs to refine uncertain reasoning steps based on these obtained manipulation results. Specifically, we construct the prompt by concatenating with manipulation results (e.g., the distance and geohash value of geo-entity) and a task instruction (i.e., ”Please refine your reasoning process”). After feeding the prompt into GPT-4, the enhanced trajectory is obtained. Detailed prompt information can be found in Appendix C.
4.3.3. Iterative Trajectory Self-refinement
Despite tool-augmented deliberation improving the geospatial computing and reasoning ability of LLMs, enhanced trajectories may not all be faithful (Hong et al., 2023). To alleviate potential error and ensure the trajectory quality (Huang et al., 2023), we refine these trajectories via an iterative self-refinement mechanism (Madaan et al., 2023). Specifically, we iterate two sequential blocks: (i) Trajectory verifier: given the trajectory, the verifier aims to provide feedback for refining the reasoning process; (ii) Trajectory updater: given the trajectory and feedback, the updater will further refine the current trajectory based on the feedback.
Trajectory Verifier. We construct a tailored prompt to ask LLMs to generate feedback. Specifically, we use a simple but effective trigger (”Judge whether all extracted triplets are correct and provide improvement suggestion”) to prompt LLMs to provide feedback. If the trajectory no longer requires modification, we let LLMs respond with ”This is a faithful trajectory”. Such a verification step lets LLMs make reflections and improve the correctness of the trajectory.
Trajectory Updater. The updater utilizes provided feedback to refine the current trajectory via prompt trigger ”Follow suggestion to refine the reasoning process”. Intuitively, the feedback may address multiple aspects (e.g., missed triplet in the RTE task or unfaithful reasoning process in the KGC task) of the unfaithful trajectories.
@@ 4.3.3. Iterative Trajectory Self-refinement
4.4. Hybrid Instruction Fine-Tuning
To construct a cost-effective UrbanKGC agent, we further utilize trajectories (generated by GPT-4) to fine-tune a smaller open-source LLM for faster inference speed and lower cost (i.e., prompting GPT-4 for UrbanKGC is expensive). Specifically, we finetune the LLM via the mixed-task instruction-tuning strategy (Zeng et al., 2023). The goal is to enhance the LLMs’ capabilities in diverse UrbanKGC tasks.
4.4.1. Mixture Training
Set the base language model as $\mathcal{M}$, and $\mathcal{P}_{\mathcal{M}}\left(y\mid x\right)$ represents the probability distribution of response $y$ given instruction $x$. We consider the trajectory set on two UrbanKGC tasks, i.e., $\mathcal{D}_{RTE}$ and $\mathcal{D}_{KGC}$. Since both the instruction and the target output are formatted in natural language, we can unify the training into an end-to-end sequence-to-sequence way. Formally, the optimization process aims to minimize the loss of language model $\mathcal{M}$ as follows:
(1) $\mathcal{L}=\mathbb{E}_{(x,y)\sim\mathcal{D}_{\text{RTE}}}\left[\log\mathcal{P}_{\mathcal{M}}(y\mid x)\right]+\mathbb{E}_{(x,y)\sim\mathcal{D}_{\text{KGC }}}\left[\log\mathcal{P}_{\mathcal{M}}(y\mid x)\right],$
@@ 4.4.1. Mixture Training
4.4.2. Training Setup
We choose the chat version of open-sourced Llama 2-7/13B (Touvron et al., 2023) as our backbone models, and fine-tune Llama using Low-Rank Adaptation (LoRA) strategy (Hu et al., 2021). Through LoRA, the training process could be done by tuning only one-thousandth of full LLM parameters.
4.5. Inference on UrbanKGC Task
Via hybrid instruction fine-tuning, the LLM can be trained to follow the instructions to finish the UrbanKGC task. After obtaining the UrbanKGent, we can prompt it to complete UrbanKGC tasks by following the pipeline shown in Figure 4.
For the relational triplet extraction task, we sequentially execute entity recognition, relation extraction, and relational triplet instruction generation. We then perform iterative self-refinement and finally output the extracted triplets. For the knowledge graph completion task, we sequentially execute KGC instruction generation, external tool augmentation, iterative self-refinement block, and finally output the completed triplets.
Table 2. The statistics of constructed UrbanKGC dataset.
Dataset NYC-Instruct NYC CHI-Instruct CHI
Records RTE 232 2,089 122 1,102
KGC 232 2,080 122 1,101
5. Experiments
We evaluate the proposed method on two real-world UrbanKGC datasets, and aim to answer the following research questions:
  • RQ1: How does the constructed UrbanKGent perform compared with existing baselines and LLM prompting paradigms on UrbanKGC tasks?
  • RQ2: How do different components (e.g., the knowledgeable instruction generation) affect the performance?
  • RQ3: How can the constructed UrbanKGent provide application service to real-world scenarios?
Table 3. The main results of relational triplet extraction (RTE) and knowledge graph completion (KGC). We report the accuracy (acc) and confidence for GPT evaluation on two datasets, and report accuracy (acc) for the Human evaluation approach. The best baseline performance is underlined.

### v2

@@ 3.1.1 Task Definition
Relational Triplet Extraction (RTE). Given the unstructured texts, this task achieves joint extraction of entities and their relations [10] which are in the form of a triplet ${\langle h,r,t\rangle}$. For instance, given the
@@
Figure 3: Quantitative performance analysis of prompting GPT-4 for UrbanKGC tasks. The result is obtained by comparing 50 GPT-4’s outputs with the human’s annotation.
urban text sentence "Columbia University is a private Ivy league research university in New York City.", this task aims to identify two entities "Columbia University" and "New York City" and their relation "locate-in", described as triplet <Columbia University, locate-in, New York City>.
Knowledge Graph Completion (KGC). Given a head entity $h$ and a tail entity $t$, this task is to predict the missing relation between them [11]. For instance, given the head entity "Columbia University" and the tail entity "Empire State Building", this task is to predict that their missing relation, e.g., "disconnected", described as triplet <Columbia University, disconnected, Empire State Building>.
3.1.2 Problem Formulation
Given the urban unstructured text data, the desired output is an UrbanKG $\mathcal{G}$. In this paper, this problem is decomposed into two sequential subtasks: (1) Relational Triplet Extraction: the first task extracts relational triplet ${\langle h,r,t\rangle}$ from the urban text data. The output of RTE task is $\mathcal{G}_{1}$ $\mathcal{E}$, $\mathcal{R}_{1}$, $\mathcal{F}_{1}$), where $\mathcal{E}$ and $\mathcal{R}_{1}$is the set of extracted entities and relations, while $\mathcal{F}_{1}$ is the set of extracted triplets. (2) Knowledge Graph Completion: for the given head entity $h$ and tail entity $t$ in $\mathcal{G}_{1}$, the second task is to predict the geospatial relationship [^7] We follow GeoLM [12] to provide five RCC relationship [30] candidates: Disconnection (DC), external connection (EC), equality (EQ), partial overlap (PO), and tangential and non-tangential proper parts (IN). Details are in Appendix A. between them. The output of this task is $\mathcal{G}_{2}$ $\mathcal{E}$, $\mathcal{R}_{2}$, $\mathcal{F}_{2}$), where $\mathcal{R}_{2}$ and $\mathcal{F}_{2}$ is the set of completed relations and triplets. By sequentially completing the above two tasks, we can obtain the constructed UrbanKG $\mathcal{G}$ = ($\mathcal{E}$, $\mathcal{R}_{1}\cup\mathcal{R}_{2}$, $\mathcal{F}_{1}\cup\mathcal{F}_{2}$).
3.2 Quantitative Task Analysis
@@
Heterogenous Relationship Understanding. The ability to understand heterogeneous relationships is ubiquitous in distilling knowledge from the massive urban corpus. For example, the text description in Figure 1 illustrates a place from spatial location, temporal time, and functional aspects. Capturing these heterogeneous semantics is important for urban knowledge distillation. We perform quantitative analysis by randomly sampling 50 urban text data and then prompt GPT-4 to complete relational triplet extraction by providing only the basic task description. As shown in Figure 3(a), we find the LLMs experience serious misjudgment (i.e., extract wrong triplets or miss the triplet) on urban spatial, temporal, and functional triplet extraction. This indicates the limited capacity of LLMs to understand heterogeneous relationships.
Geospatial Computing and Reasoning. Geospatial computing and reasoning techniques are widely used in many previous UrbanKG studies [8, 24] for urban geospatial relation extraction. In recent works [23, 31], the geospatial skills of LLMs have also been demonstrated to lack geospatial awareness and reasoning ability [22]. To identify potential limitations, we quantitatively investigate how LLMs can perform on geospatial relation completion tasks. Specifically, we construct 100 head and tail entity pairs, covering five geospatial relations in the KGC task, and then prompt GPT-4 to predict with basic task description and geospatial relation candidates. As shown in Figure 3(b), we find that GPT-4 performs poorly on five geospatial relation completion. This further validates the disability of LLMs in geospatial computing and reasoning.
4 UrbanKGC Agent Construction
@@
4.1 Overview
The overall pipeline of the UrbanKGent framework is illustrated in Figure 4. (1) Knowledgeable Instruction Generation consists of the heterogeneity-aware and geospatial-infused instruction generation modules for aligning LLMs to UrbanKGC tasks. (2) Tool-augmented Iterative Trajectory Refinement proposes geospatial tool interface invocation and iterative self-refinement mechanisms to enhance and refine generated trajectory. (3) Hybrid Instruction Fine-tuning fine-tune LLMs based on the refined trajectories for cost-effectively completing diverse UrbanKGC tasks.
4.2 Knowledgeable Instruction Generation
@@
Heterogeneity-aware Instruction Generation for Relational Triplet Extraction. As discussed in Section 3, the urban text contains diverse heterogeneous relationships, thus we consider multiple views with both urban entity and relation definition for relational triplet extraction. In particular, we construct a multi-view instruction template for the urban relational triplet extraction, including spatial view, temporal view, and functional view. Each view is a multi-turn question-answer dialog [32] consisting of entity recognition, relation extraction, and triplet extraction module.
@@
For the spatial view, we devise a two-turn dialog to align LLMs for spatial triplet extraction. In the first turn, we inject spatial entity and relation definition into the instruction template to guide LLMs to understand spatial characteristics and then extract potential spatial entities (e.g., University) and relations types (e.g., locate-in). In the second turn, the extracted types are explicitly fed into the instruction template for spatial triplet extraction. Intuitively, the spatial view allocates dedicated urban knowledge for LLMs to extract urban spatial relationships. Similarly, we construct the temporal view and functional view for corresponding temporal and functional triplet extraction, independently.
@@
Geospatial-infused Instruction Generation for Knowledge Graph Completion. Despite heterogeneity-aware instruction enabling LLMs to extract urban triplets from various perspectives, the geospatial relationship between geospatial entities cannot be directly extracted. Therefore, we introduce a geospatial-infused instruction generation module to guide LLMs to complete missing geospatial relationships.
@@
First, we incorporate the geometry information (i.e., the latitude and longitude) of geo-entities into instruction, so that the LLMs can utilize these geospatial values for relation inference. Second, we add the geospatial relationship definition to the instruction to guide LLMs in understanding the geospatial relationship definition. Intuitively, LLMs can refer to geospatial knowledge and make practical solutions for the KGC task. We provide the detailed instruction template in the Appendix B.
4.3 Tool-augmented Iterative Trajectory Refinement
Figure 4: An overview of UrbanKGent Construction.
4.3.1 Trajectory Generation
@@
With the initial UrbanKGC instructions constructed, the following step is to generate reasoning trajectories [33], which will be used to fine-tune LLMs tailored to UrbanKGC task. Specifically, we follow FireAct [34] and use Chain-of-Thought (CoT) [35], a gradient-free technique, to prompt GPT-4 (i.e., add prompt trigger "Let’s think step by step" at the end of RTE and KGC instructions template) to generate the reasoning trajectories for UrbanKGC tasks.
@@
The generated CoT trajectories could provide a step-by-step reasoning solution for UrbanKGC tasks. Nevertheless, the complex geospatial relationships cannot be easily extracted as discussed in Section 3 and recent geospatial reasoning works [31, 23, 22]. Therefore, we introduce a tool invocation module to guide LLMs to invoke tailored external geospatial tools [36] to enhance their geospatial computing and reasoning capacity for UrbanKGC tasks.
4.3.2 Tool Invocation for Trajectory Augmentation
@@
Tool Invocation. First, we construct a geospatial reasoning toolkit (e.g., distance calculation, eight interfaces in total shown in Table 6) by prompting GPT-4 for self-programming. Then, we construct tailored prompts to guide LLMs to invoke these interfaces. Specifically, the prompt is concatenated with an illustrative description of the function of each geospatial tool and a task instruction (i.e., "Which types of tool interface you need"). Intuitively, the external tool allocates calculation results for LLMs to infer missing geospatial relation. The toolkit description can be found in Appendix C.
Trajectory Deliberation. After manipulation with external tools, we prompt LLMs to refine uncertain reasoning steps based on these obtained manipulation results. Specifically, we construct the prompt by concatenating with manipulation results (e.g., the distance and geohash value of geo-entity) and a task instruction (i.e., "Please refine your reasoning process"). After feeding the prompt into GPT-4, the enhanced trajectory is obtained. Detailed prompt information can be found in Appendix C.
4.3.3 Iterative Trajectory Self-refinement
Despite tool-augmented deliberation improving the geospatial computing and reasoning ability of LLMs, enhanced trajectories may not all be faithful [37]. To alleviate potential error and ensure the trajectory quality [38], we refine these trajectories via an iterative self-refinement mechanism [39]. Specifically, we iterate two sequential blocks: (i) Trajectory verifier: given the trajectory, the verifier aims to provide feedback for refining the reasoning process; (ii) Trajectory updater: given the trajectory and feedback, the updater will further refine the current trajectory based on the feedback.
Trajectory Verifier. We construct a tailored prompt to ask LLMs to generate feedback. Specifically, we use a simple but effective trigger ("Judge whether all extracted triplets are correct and provide improvement suggestion") to prompt LLMs to provide feedback. If the trajectory no longer requires modification, we let LLMs respond with "This is a faithful trajectory". Such a verification step lets LLMs make reflections and improve the correctness of the trajectory.
Trajectory Updater. The updater utilizes provided feedback to refine the current trajectory via prompt trigger "Follow suggestion to refine the reasoning process". Intuitively, the feedback may address multiple aspects (e.g., missed triplet in the RTE task or unfaithful reasoning process in the KGC task) of the unfaithful trajectories.
@@ 4.3.3 Iterative Trajectory Self-refinement
4.4 Hybrid Instruction Fine-Tuning
To construct a cost-effective UrbanKGC agent, we further utilize trajectories (generated by GPT-4) to fine-tune a smaller open-source LLM for faster inference speed and lower cost (i.e., prompting GPT-4 for UrbanKGC is expensive). Specifically, we finetune the LLM via the mixed-task instruction-tuning strategy [33]. The goal is to enhance the LLMs’ capabilities in diverse UrbanKGC tasks.
Mixture Training. Set the base language model as $\mathcal{M}$, and $\mathcal{P}_{\mathcal{M}}\left(y\mid x\right)$ represents the probability distribution of response $y$ given instruction $x$. We consider the trajectory set on two UrbanKGC tasks, i.e., $\mathcal{D}_{RTE}$ and $\mathcal{D}_{KGC}$. Since both the instruction and the target output are formatted in natural language, we can unify the training into an end-to-end sequence-to-sequence way. Formally, the optimization process aims to minimize the loss of language model $\mathcal{M}$ as follows:
$\mathcal{L}=\mathbb{E}_{(x,y)\sim\mathcal{D}_{\text{RTE}}}\left[\log\mathcal{P}_{\mathcal{M}}(y\mid x)\right]+\mathbb{E}_{(x,y)\sim\mathcal{D}_{\text{KGC }}}\left[\log\mathcal{P}_{\mathcal{M}}(y\mid x)\right],$ (1)
@@ 4.4 Hybrid Instruction Fine-Tuning
Training Setup. We choose the chat version of open-sourced Llama 2-7/13B and Llama-3-8B as our backbone models, and fine-tune Llama using LoRA strategy [26].
4.5 Inference on UrbanKGC Task
Via hybrid instruction fine-tuning, the obtained LLM UrbanKGent can be trained to follow the instructions to finish the UrbanKGC task. We prompt UrbanKGent to complete UrbanKGC tasks by following the pipeline shown in Figure 4. For the RTE task, we sequentially execute entity recognition, relation extraction, and relational triplet instruction generation, iterative self-refinement and output the extracted triplets. For the KGC task, we sequentially execute KGC instruction generation, external tool augmentation, iterative self-refinement block, and finally output the completed triplets.
Table 2: The statistics of constructed UrbanKGC dataset.
Dataset NYC-Instruct NYC NYC-Large CHI-Instruct CHI CHI-Large
Records RTE 232 2,089 40,480 122 1,102 28,868
KGC 232 2,080 33,534 122 1,101 28,607
5 Experiments
5.1 Experimental Settings
Dataset. In this work, two sequential tasks (i.e., RTE and KGC) of UrbanKGC are within an open-world setting (i.e., no predefined ontology) [40, 41]. We construct the RTE and KGC datasets of NYC and CHI by sampling uniformly from five raw data in Table 1, respectively. As shown in Table 2, we first construct two small datasets (i.e., NYC-Instruct and CHI-Instruct) for instruction fine-tuning and two middle datasets (i.e., NYC and CHI) to validate the performance of the constructed UrbanKGC agent. The remaining data works as the large-scale UrbanKGC dataset (i.e., NYC-Large and CHI-Large) in real-world scenarios shown in Table 5. The three types of datasets are non-overlapping to prevent data leakage. More dataset construction details are in Appendix A.
Table 3: The main results of relational triplet extraction (RTE) and knowledge graph completion (KGC). We report the accuracy (acc) and confidence for GPT evaluation on two datasets, and report accuracy (acc) for the Human evaluation approach. The best baseline performance is underlined.